OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + 2y = 0 2x + 4y = 0

OpenStudy (anonymous):

Use elimination, x and y both cancel out, so it's 0=0. That's infinite.

OpenStudy (nbarrera):

Multiply both sides of the first equation by 2.

OpenStudy (anonymous):

2x+4y=0 @nbarrera

OpenStudy (anonymous):

Noo, multiply them by -2. Then add the top and botton equations.

OpenStudy (anonymous):

oh okay so then both sides equal zero @logancoop

OpenStudy (anonymous):

Yep, and since 0=0 it's infinite.

OpenStudy (anonymous):

@madison.bush if multiplying or dividing any of equation, you get second one, then solutions are infinite.

OpenStudy (anonymous):

Like, if I multiply 2 with your first equation, we will get second equation.

OpenStudy (anonymous):

infinite is not one of my answer choices. they are one solution no solution coincident

OpenStudy (anonymous):

\(2(x+2y = 0)\) \(\implies 2x + 4y = 0\), this is same as your second equation.

OpenStudy (anonymous):

Coincident = Infinite number Solutions.

OpenStudy (anonymous):

Coincident means the two lines are fully overlapping over each other. :)

OpenStudy (anonymous):

OHHHH THANKS!

OpenStudy (anonymous):